# Entropy conservative and entropy stable solid wall boundary conditions for the resistive magnetohydrodynamic equations

### v1

@@ 5.3 Fluid flow in microchannel
At the inlet, we impose an uniform normal velocity, an external magnetic field $\vec{B}^{0}$ and zero normal gradients. The inlet velocity magnitude is defined such that the inlet mass flow rate at time step $n+1$ equals the outlet mass flow rate at the previous time step $n$. The initial mass flow rate is zero. With the notation from Sections 4.1 and 4.2, the inlet manufacture state and gradients are defined as
@@ 5.4 Self-field magnetoplasmadynamic thruster
[…]
$\mathbf{v}_{(+)}^{\text{outlet}}=\left[\rho_{(-)}\left(\frac{p^{\text{out}}}{p_{(-)}}\right)^{\frac{1}{\gamma}},\vec{V}^{\text{out}},T_{(-)}\left(\frac{p^{\text{out}}}{p_{(-)}}\right)^{\frac{\gamma-1}{\gamma}},\vec{0},0\right]^{T},\quad\theta_{(+)}^{\text{inlet}}=\theta_{(-)}-(\theta_{(-)}\cdot\vec{n})\,\vec{n},$ (5.22)
where $\vec{V}^{\text{out}}=\vec{v}_{(-)}-2\min{(\vec{n}\cdot\vec{v}_{(-)},0)}\,\vec{n}$ is the velocity of the gas with a simple limiter to prevent backflow if $\vec{n}\cdot v_{(-)}<0$. Density $\rho_{(+)}$ and temperature $T_{(+)}$ are defined such that the pressure $p_{(+)}=p^{\text{out}}$ and the specific entropy equals that of the internal state, that is $s_{(+)}=s_{(-)}$.
@@
[…]
Due to the axial symmetry of the solution, the electric current has a stream function $\Phi=rB_{\theta}$ shown in Figure 9. Electric current streamlines lie on iso-surfaces of $\Phi$. Consistently with the continuous analysis, all the electric current streamlines originate at the cathode surface and end at the anode surface. The gas tends to convect electric current streamlines downstream of the thruster chamber, while the electrical resistivity keeps such spread within the nearby region.
@@ 6 Conclusions
We have used entropy stability and the summation-by-parts framework to derive entropy conservative and entropy stable wall boundary conditions for the resistive magnetohydrodynamic equations in the presence of an adiabatic wall or a wall with a prescribed heat entropy flow, addressing three scenarios: electrically insulating walls, thin walls with finite conductivity, and perfectly conducting walls. A point-wise entropy-stable numerical procedure has been presented for weakly enforcing these solid wall boundary conditions at the semi-discrete level, combining a penalty flux approach with a simultaneous approximation-term technique. The resulting semi-discrete operator mimics exactly the behavior at the continuous level, and the proposed non-linear boundary treatment provides a mechanism for ensuring non-linear stability in the $L_{2}$ norm of the continuous and semi-discretized resistive magnetohydrodynamic equations.
@@ Appendix A Entropy stable two-point flux function
Appendix B Jacobian of the transformation from entropy to conservative variables
The Jacobian of the transformation matrix from conservative variables, $\mathbf{u}$, to entropy variables $\mathbf{w}$, for GLM-MHD system, i.e., $\mathcal{H}={\partial\mathbf{u}}/{\partial\mathbf{w}}$ is defined in [19]. In the following, we reproduce its expression. Our definition has an additional $1/R$ scaling factor related to our choice of entropy function, $S$, in (2.14) and (2.15).
Consider a state of the GLM-MHD system defined in primitive variables as $\mathbf{v}=(\rho,v_{1},v_{2},v_{3},T,B_{1},B_{2},B_{3},\psi)$. Then,
$\mathcal{H}$ $=\frac{1}{R}\begin{bmatrix}\rho&\rho v_{1}&\rho v_{2}&\rho v_{3}&\mathcal{E}&0&0&0&0\\ \rho v_{1}&\rho v_{1}^{2}+p&\rho v_{1}v_{2}&\rho v_{1}v_{3}&(\mathcal{E}+p)v_{1}&0&0&0&0\\ \rho v_{2}&\rho v_{2}v_{1}&\rho v_{2}^{2}+p&\rho v_{2}v_{3}&(\mathcal{E}+p)v_{2}&0&0&0&0\\ \rho v_{3}&\rho v_{3}v_{1}&\rho v_{3}v_{2}&\rho v_{3}^{2}+p&(\mathcal{E}+p)v_{3}&0&0&0&0\\ \mathcal{E}&(\mathcal{E}+p)v_{1}&(\mathcal{E}+p)v_{2}&(\mathcal{E}+p)v_{3}&\mathcal{H}_{5,5}&\tau B_{1}&\tau B_{2}&\tau B_{3}&\tau\psi\\ 0&0&0&0&\tau B_{1}&\tau\mu_{0}&0&0&0\\ 0&0&0&0&\tau B_{2}&0&\tau\mu_{0}&0&0\\ 0&0&0&0&\tau B_{3}&0&0&\tau\mu_{0}&0\\ 0&0&0&0&\tau\psi&0&0&0&\tau\mu_{0}\\ \end{bmatrix},$ (B.1)
$\tau$ $=\frac{p}{\rho}=RT$ (B.2)
$\mathcal{E}$ $=\frac{p}{\gamma-1}+\rho\frac{|\!|\vec{v}|\!|^{2}}{2},$ (B.3)
$\mathcal{H}_{5,5}$ $=\frac{1}{\rho}\left(\frac{p^{2}}{\gamma-1}+\mathcal{E}^{2}\right)+p|\!|\vec{v}|\!|^{2}+\frac{\tau}{\mu_{0}}\left(|\!|\vec{B}|\!|^{2}+\psi^{2}\right).$ (B.4)
@@ Appendix D Analytical solution for MHD pipe flow
In this section, we present an analytical solution to the flow of a conducting fluid in a straight pipe of circular cross-section under uniform traverse magnetic field. This analytical solution is derived in [59] and used in section 5.1 to validate the accuracy of the solid wall boundary conditions we propose in this work. An uniform external force in the axial direction causes the fluid motion, and the flow is assumed incompressible.
@@ Appendix F Magnetic field of a wire with rectangular cross-section
Here, we reproduce an analytical expression for the magnetic field induced by an uniform electric current flowing through a conducting medium of infinite length with rectangular cross section of width $w$ and thickness $h$ [61]. This analytical solution can be used to approximate the solution of a finite domain of length $L$ with $w\ll L$ and $h\ll L$. In Cartesian coordinates, with the current of density $i$ flowing in positive $y$ direction through the wire of rectangular cross section $(x,z)\in[-w/2,w/2]\times[-h,0]$, the components of the induced magnetic field $\vec{B}$ are
@@ Appendix G Magnetic field of a circular current loop
$B_{r}$ $=B_{0}\frac{\gamma}{\pi\sqrt{Q}}\left(E(k)\frac{1+\alpha^{2}+\beta^{2}}{Q-4\alpha}-K(k)\right),\quad B_{\theta}=0,$

### v2

@@ 5.3 Fluid flow in microchannel
At the inlet, we impose a uniform normal velocity profile, an external magnetic field $\vec{B}^{0}$ and zero normal gradients. The inlet velocity magnitude is defined such that the inlet mass flow rate at time step $n+1$ equals the outlet mass flow rate at the previous time step $n$. The initial mass flow rate is zero. With the notation from Sections 4.1 and 4.2, the inlet manufacture state and gradients are defined as
@@ 5.4 Self-field magnetoplasmadynamic thruster
The computational domain has a cylindrical shape with a diameter and length equal to six cathode lengths. It is large enough for the far-field outlet boundaries to not noticeably to affect the solution, and such that all significant physical phenomena are captured within the region of interest close to the thruster. Outlet boundary conditions include zero magnetic field and external pressure $p^{\text{out}}=0.1p^{0}$, approximating a vacuum condition. In the notation of Section 4.1 and 4.2, outlet boundary conditions are
$\mathbf{v}_{(+)}^{\text{outlet}}=\left[\rho_{(-)},\vec{V}^{\text{out}},T_{(-)}\left(\frac{p^{\text{out}}}{p_{(-)}}\right),\vec{0},0\right]^{T},\quad\theta_{(+)}^{\text{inlet}}=\theta_{(-)}-(\theta_{(-)}\cdot\vec{n})\,\vec{n},$ (5.22)
where $\vec{V}^{\text{out}}=\vec{v}_{(-)}-2\min{(\vec{n}\cdot\vec{v}_{(-)},0)}\,\vec{n}$ is the velocity of the gas with a simple limiter to prevent backflow if $\vec{n}\cdot v_{(-)}<0$. The density $\rho_{(+)}$ and the temperature $T_{(+)}$ are defined such that the pressure satisfies $p_{(+)}=p^{\text{out}}$.
@@
Figure 9: Princeton full-scale benchmark thruster: steady state solution for $\operatorname{Ma}=0.4$, $\operatorname{M_{m}}=0.5$.
In agreement with experimental observations and previous numerical simulations from [64], the plasma accelerates as it passes through the thruster chamber, as shown in Figure 9. Due to the axial symmetry of the solution, the electric current has a stream function $\Phi=rB_{\theta}$ shown in Figure 9. Electric current streamlines lie on iso-surfaces of $\Phi$. Consistently with the continuous analysis, the electric current streamlines which originate at the cathode surface end at the anode surface. The gas tends to convect electric current streamlines downstream of the thruster chamber, while the electrical resistivity keeps such spread within the nearby region.
@@ 6 Conclusions
We have used entropy stability and the summation-by-parts framework to derive entropy conservative and entropy stable wall boundary conditions for the resistive magnetohydrodynamic equations in the presence of an adiabatic wall or a wall with a prescribed heat entropy flow, addressing three scenarios: electrically insulating walls, thin walls with finite conductivity, and perfectly conducting walls. A point-wise entropy-stable numerical procedure has been presented for weakly enforcing these solid wall boundary conditions at the semi-discrete level, combining a penalty flux approach with a simultaneous approximation-term technique. The resulting semi-discrete operator mimics the behavior at the continuous level, and the proposed non-linear boundary treatment provides a mechanism for ensuring non-linear stability in the $L_{2}$ norm of the continuous and semi-discretized resistive magnetohydrodynamic equations.
@@ Appendix A Entropy stable two-point flux function
Appendix B Numerical entropy Jacobian used in the definition of LLF dissipation
The numerical entropy Jacobian $\mathcal{H}$ designed for GLM-MHD system in [19] completes the definition of LLF operator (4.3). In the following, we reproduce the expression for $\mathcal{H}$ reusing definitions from A.
Consider a state of the GLM-MHD system defined in primitive variables as $\mathbf{v}=(\rho,u,v,w,T,B_{1},B_{2},B_{3},\psi)$. Then,
$\mathcal{H}=\mathcal{H}(\mathbf{v}_{i},\mathbf{v}_{j})=$
$\frac{1}{R}\begin{bmatrix}{\rho}^{\text{ln}}&{\rho}^{\text{ln}}\{\!\!\{u\}\!\!\}&{\rho}^{\text{ln}}\{\!\!\{v\}\!\!\}&{\rho}^{\text{ln}}\{\!\!\{w\}\!\!\}&\overline{E}&0&0&0&0\\ {\rho}^{\text{ln}}\{\!\!\{u\}\!\!\}&{\rho}^{\text{ln}}\{\!\!\{u\}\!\!\}^{2}+\overline{p}&{\rho}^{\text{ln}}\{\!\!\{u\}\!\!\}\{\!\!\{v\}\!\!\}&{\rho}^{\text{ln}}\{\!\!\{u\}\!\!\}\{\!\!\{w\}\!\!\}&(\overline{E}+\overline{p})\{\!\!\{u\}\!\!\}&0&0&0&0\\ {\rho}^{\text{ln}}\{\!\!\{v\}\!\!\}&{\rho}^{\text{ln}}\{\!\!\{v\}\!\!\}\{\!\!\{u\}\!\!\}&{\rho}^{\text{ln}}\{\!\!\{v\}\!\!\}^{2}+\overline{p}&{\rho}^{\text{ln}}\{\!\!\{v\}\!\!\}\{\!\!\{w\}\!\!\}&(\overline{E}+\overline{p})\{\!\!\{v\}\!\!\}&0&0&0&0\\ {\rho}^{\text{ln}}\{\!\!\{w\}\!\!\}&{\rho}^{\text{ln}}\{\!\!\{w\}\!\!\}\{\!\!\{u\}\!\!\}&{\rho}^{\text{ln}}\{\!\!\{w\}\!\!\}\{\!\!\{v\}\!\!\}&{\rho}^{\text{ln}}\{\!\!\{w\}\!\!\}^{2}+\overline{p}&(\overline{E}+\overline{p})\{\!\!\{w\}\!\!\}&0&0&0&0\\ \overline{E}&(\overline{E}+\overline{p})\{\!\!\{u\}\!\!\}&(\overline{E}+\overline{p})\{\!\!\{v\}\!\!\}&(\overline{E}+\overline{p})\{\!\!\{w\}\!\!\}&R\mathcal{H}_{5,5}&\tau\{\!\!\{B_{1}\}\!\!\}&\tau\{\!\!\{B_{2}\}\!\!\}&\tau\{\!\!\{B_{3}\}\!\!\}&\tau\{\!\!\{\psi\}\!\!\}\\ 0&0&0&0&\tau\{\!\!\{B_{1}\}\!\!\}&\tau\mu_{0}&0&0&0\\ 0&0&0&0&\tau\{\!\!\{B_{2}\}\!\!\}&0&\tau\mu_{0}&0&0\\ 0&0&0&0&\tau\{\!\!\{B_{3}\}\!\!\}&0&0&\tau\mu_{0}&0\\ 0&0&0&0&\tau\{\!\!\{\psi\}\!\!\}&0&0&0&\tau\mu_{0}\\ \end{bmatrix},$
$\overline{E}=\frac{1}{\gamma-1}\frac{{\rho}^{\text{ln}}}{{\beta}^{\text{ln}}}+{\rho}^{\text{ln}}\left(\{\!\!\{u\}\!\!\}^{2}+\{\!\!\{v\}\!\!\}^{2}+\{\!\!\{w\}\!\!\}^{2}\right)-{\rho}^{\text{ln}}\frac{\{\!\!\{u^{2}+v^{2}+w^{2}\}\!\!\}}{2},\quad\tau={\{\!\!\{\beta\}\!\!\}}^{-1},$
$R\mathcal{H}_{5,5}=\frac{\{\!\!\{u\}\!\!\}^{2}+\{\!\!\{v\}\!\!\}^{2}+\{\!\!\{w\}\!\!\}^{2}}{2}\left(\overline{E}+\overline{p}\right)+\tau\left(\overline{E}+\frac{\{\!\!\{B_{1}\}\!\!\}^{2}+\{\!\!\{B_{2}\}\!\!\}^{2}+\{\!\!\{B_{3}\}\!\!\}^{2}}{\mu_{0}}+\frac{\{\!\!\{\psi\}\!\!\}^{2}}{\mu_{0}}\right).$
@@ Appendix D Analytical solution for MHD pipe flow
In this section, we present an analytical solution to the flow of a conducting fluid in a straight pipe of circular cross-section under uniform traverse magnetic field. This analytical solution is derived in [58] and used in section 5.1 to validate the accuracy of the solid wall boundary conditions we propose in this work. A uniform external force in the axial direction causes the fluid motion, and the flow is assumed incompressible.
@@ Appendix F Magnetic field of a wire with rectangular cross-section
Here, we reproduce an analytical expression for the magnetic field induced by a uniform electric current flowing through a conducting medium of infinite length with rectangular cross section of width $w$ and thickness $h$ [60]. This analytical solution can be used to approximate the solution of a finite domain of length $L$ with $w\ll L$ and $h\ll L$. In Cartesian coordinates $(x,y,z)$, with the current of density $i$ flowing in positive $y$ direction through the wire of rectangular cross section $(x,z)\in[-w/2,w/2]\times[-h,0]$, the components of the induced magnetic field $\vec{B}$ are
@@ Appendix G Magnetic field of a circular current loop
$B_{r}$ $=B_{0}\frac{\gamma}{\pi\sqrt{Q}}\left(E(k)\frac{1+\alpha^{2}+\beta^{2}}{Q-4\alpha}-K(k)\right),$
$\quad B_{\theta}$ $=0,$